OpenStudy (anonymous):

when working with polynomials, if the problem says "EVALUATE" does tht means solve or just state the problem ?

OpenStudy (oaktree):

Simplify. Find the final value.

OpenStudy (mertsj):

It means replace the variables with the specified numbers and see what it turns out to be.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what mertzj said evaluate \(4x^2+3x+1\) at \(x=10\) gives \(4\times 10^2+3\times 10+1=431\)